OpenStudy (anonymous):

Math check! Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car.

OpenStudy (anonymous):

Part A: Write functions to represent Cory and Roger's collections throughout the years. Answer: Cory: y = 15(1.2)^ x Roger y = x + 40

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years?

OpenStudy (anonymous):

Rodger will have 46 cars while Cory will have 55

OpenStudy (anonymous):

the 46 is certainly right need a calculator for the other one

OpenStudy (anonymous):

that is not what i get be careful how you enter it

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

Wait... would it be 52?

OpenStudy (anonymous):

this is what i got http://www.wolframalpha.com/input/?i=15%281.2%29^6

OpenStudy (anonymous):

okay so it would have been 44

OpenStudy (anonymous):

depends on how you round, but i guess so if you round down

OpenStudy (anonymous):

Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically.

OpenStudy (anonymous):

cory: y = 15 * 1.2^x roger: y = 40 + x the functions intersect when x = 6.16599

OpenStudy (anonymous):

this requires a calculator, but when you approximate i guess they may want you to round

OpenStudy (anonymous):

i get the same thing you do

OpenStudy (anonymous):

yeah, Ive been having a hard time with this part

OpenStudy (anonymous):

Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically.

OpenStudy (anonymous):

that is the same as what you just did

OpenStudy (anonymous):

approximately 6 i would say

OpenStudy (anonymous):

After 6 years Rodger will have 40 cars while cory will have 44?

OpenStudy (anonymous):

46

OpenStudy (anonymous):

after 6 years roger will have 46 right?

OpenStudy (anonymous):

yeah, i have to find when they would have the same amount of cars

OpenStudy (anonymous):

it will not be a whole number of years at 6 roger has 46 and cory has 44.8 or so but after 7 years roger will have 47 and cory will have \(15(1.2)^7=53.74\) cars so year 6 is the closest

OpenStudy (anonymous):

Okay, thank you! I've been stuck on this all day.